Plan for E/D Nights UT Dec 28-Dec 31, Jan02

UT Dec 28, 29, 30, 31 - 2nd half of the night, Jan 02 - full night

Tasks to be done:

Engineering Tasks
  • DONE (done 20201227 UT) SX M3 alignment after reinstallation ( CRITICAL )
    • SX M3 pupil wobble measurement and M3 alignment (2 hours, any seeing, modest clouds)
    • Commissioning/BGTechProcedures#Procedure_for_Measuring_Pupil_Wobble_and_Adjusting_M3_with_Off_45Axis_Guider
      • Take hint to use 5 or 6 positions per rotation rather than 4.
  • DONE Deploy new TCS 2021A - high priority - (daytime testing completed 20201224)
  • SOUL AO Checkout following problems of 20201220 UT ( CRITICAL , must follow above tasks)
    • scripts /home/lbto/Service_Observing/D_E_Time_UT_Dec28_Jan_02/AO_Testing/
    • DONE Bin 3 test camerea limit with multiple rotator angles
    • Bin 1-4 testing
    • Full AO using N375 camera (ESM fields) - scripts in /home/lbto/Service_Observing/D_E_Time_UT_Dec28_Jan_02/ESM_Testing
  • LBC Technical Work - Medium Priority
    - O.Kuhn running Technical Work
    • focus offset checks - stable seeing (1" or better on DIMM) - 1.5hr
      • scripts in lbto@robs:~/Service_Observing_D_E_Time_UT_Dec28_Jan02/LBC_R_focus
      • check LBCR R focus, mainly.
        • Focus/collimate, then run superfoc at V+r, followed by a superfoc at [Us,B,g]+R, followed again by superfoc at V+r. Repeat 3x.
    • Gain/Linearity checks - 1hr (photometric conditions, but poor seeing OK)
      • scripts in lbto@robs:~/Service_Observing_D_E_Time_UT_Dec28_Jan02/LBC_Linearity
      • observe a star field with successively longer exposure times, so that a set of stars will have counts that range from 1000 ADU to saturation.
      • several star fields will be chosen, with published photometry (SDSS or other) so we can not only examine the linearity but check the sensitivity.
    • TMS: Run TMS in active mode while collecting a series of pupil images (dfoc = -0.8, Z4/Z11 as sent by fpia) - 2hr any conditions
  • ESM checkout - high priority
    • Scripts: /home/lbto/Service_Observing/D_E_Time_UT_Dec28_Jan_02/ESM_Testing
    • On-sky checks Bin 2 400Hz (FLAO config) vs Bin 2 100Hz (SOUL config) - 1-2 hrs
      • Collect ESM data - <1" seeing - stable - 1hr minmum
    • Collect Palomar 2 SOUL-ESM data - clear and <1" seeing
      • PA0 and original files are in directory
      • Please collect Phot standards before and after observations
      • This is to compare with similar data taken with FLAO
    • FULL AO using N375 camera for M52 or M67 with bright and faint AO Ref stars
    • DONE NGC 3921 SOUL ESM for comparison with FLAO (2nd half)
  • MODS check - Medium Priority
    • Script Locations: /home/lbto/Service_Observing/D_E_Time_UT_Dec28_Jan_02/MODS_Checkout
    • Check rotator angle duirng a MODS MOS acquisition (Stone fields) - High Priority - 20min
    • Fields:
      • StoneA - sets 10pm
      • Stone D1f - midnight-2am
      • Stone E1f - midnight-5am
      • Stone I - 3am-twilight
      • MMS files in mms/ sub-directory
    • AGW confirmatory transform (1 hour)
  • LBTI Technical Work (1/2 night)
    • Engineering tests at start of night - spillover a couple of hour into the 2nd half
      • best weather preferred - otherwise no preference
    • Gopt and NCPA test (DIMM ~1" or better, thin clouds at most)
    • Calibrating ACE tables (DIMM ~1.5" or better, clear-ish)
    • measuring contrast of vAPP (DIMM ~1" or better, some clouds OK)
    • testing the AGPM centering for MPIA time (DIMM ~1.2" or better, thin clouds at most)

Science Programs - D-Time
  • UDS z~10 galaxy - L1/ESM and L2/SL - UT Jan 02 - 1st half - 3hrs
    • requires DIMM 1" or better, clear skies or thin stable clouds
    • Scripts: /home/lbto/Service_Observing/D_E_Time_UT_Dec28_Jan_02/UDS_z10_ESM_Science
  • HD 33632 observations (first-half night target)
    • 1.5 hours with LUCI-1 AO (Steve E's suggestion: (DIMM ~1.2" or better, thin clouds at most)
    • 3 hours with LBTI for an ALES L band spectrum (DIMM ~1.2" or better, clear, dry)
    • 2 hours with LBTI for M band photometry (DIMM ~1.2" or better, thin clouds at most)
    • best done in three separate nights to observe each around transit, LUCI + LBTI M band might just be done during one night, but we'll lose data quality from reduced field rotation.
    • First night could be done UT Dec 29

Poor Weather/Poor Seeing E/D Programs
  • ESM data collection M34 (3 filters) , M52 (3 filters and bright and faint AO Ref star), M67 (3 filters and bright and faint AO Ref star) - if you can close the loop take the data
    • this is different than ESM tests above - this is sytematic data colleciton - above is checking on paramters for optimal use
    • Scripts: /home/lbto/Service_Observing/D_E_Time_UT_Dec28_Jan_02/ESM_Testing
  • ICRF3 Spectroscopy (Rothberg)
    • Scripts:/home/lbto/Service_Observing/D_E_Time_UT_Dec28_Jan_02/Poor_Weather/ICRF3_Rothberg_2020B
    • Observe telluric first
    • Double Check high Dec objects to make sure tellurc/science are close spatially
  • LUCI Photometric Standards
    • Scripts: /home/lbto/Service_Observing/D_E_Time_UT_Dec28_Jan_02/LUCI_Phot_ZP
    • Can be done in ANY SEEING - but it must be clear
    • Select stars to use based on visibility - OT file is in directory, export the scripts as needed
      • Please observe the same star at least two times at different airmasses
Closed Dome Tasks
  • DONE INAF Darks
    • /home/inaf//ScienceRuns/2019/LUCI/
      • DONE Nardini_4/cal/darks_nov2020_bino.xml
      • DONE Davanzo_DDT-2019B-9/calib/darks_luci1_savecube.xml
  • DONE AZ Darks
    • /Service_Observing/WASP12_transit/L1_25darks.xml
  • Calibration Frames
    • DONE ESM Imaging N375 J,H,K,Ks - bino
    • G210 z,J,H,K for ICR5 target - poor weather
    • G200 zJ, HK G220 for ICRF3 - poor weather
    • AO imaging - J,H,K,Ks N30 - LUCI-1
  • ESM Closed Dome tests
    • DONE Update TTM table
    • DONE Test TTM table w/WFS
    • DONE Obtain LUCI images w/ARGOS cal unit with new TTM table
      • DONE could sample more Bins/mags
Night to Night
  • UT Dec 28 - Night Plan - opened short time for LBTI only, closed dome Eng 2nd half
  • UT Dec 29 Prelim
    • Rolled over UT Dec 28 plan - never opened, closed dome Eng
  • UT Dec 30 Prelim
    • Rolled Over UT Dec 28 plan - never opened - ice and snow
  • UT Dec 31 Prelim
    • Rolled over UT Dec 28 plan


-- BarryRothberg - 09 Dec 2020
Topic revision: r32 - 01 Jan 2021, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback